1 Binomiais

Para escrever binomiais, usamos \binom{n}{p}. Por exemplo:

\binom{a + b}{x^{2}}

2 Divisão

\frac{dividendo}{divisor}. Exemplo:

f(x) = \frac{x^{2} + 2x}{\sqrt[4]{x}}

3 Somatório

\sum. Exemplo:

f(x) = ∑_{n=0}^{10} 2x + ∑_{n=0}^{5} x^{2}